# {refsection}

### v1

@@ Introduction
To complement the mathematical model and to further investigate the mechanical origins of the resistance of division-formed colonies to intense shear, we performed rheology tests on the cyanobacterial colonies (Materials and Methods). Simple and oscillatory shear measurements revealed elasto-viscoplastic mechanics [balmforth2014yielding, bonn2017yield] of the concentrated aggregate with a yield stress of $\tau_{y}=4.3\pm 0.3\text{ Pa}$. The low volume fraction of cells in aggregate ($\sim 12$) indicates that the mechanical properties are mainly dictated by the EPS layer. Equating the average hydrodynamic shear stress to the yield stress results in a critical colony size of $l^{*}_{2}=2.8$ (Eq. 10 in Materials and Methods). The eroded fragments observed in the fragmentation experiments (Fig. 2E) are mostly smaller than this critical size. This suggests that large colonies will suffer erosion when the local hydrodynamic stress exceeds the yield stress of the EPS layer. Although the fracture behavior of colonies in suspension is expected to be different from a concentrated aggregate under shear rheology, the latter provides a reasonable and convenient estimate of the material properties of the colonies. Related to the critical size of the colony and their mechanical properties is the exponent $q_{2}$ (see equation 4). The best fit of the results in Fig. 2 provided $q_{2}=4.1$. This value lies in the upper limit of previous results for colloidal aggregates [flesch1999laminar, baebler2008modelling]. Note that for colloids, $q$ is related to the stress concentration around cracks in the fractal structure of the aggregate [zaccone2009breakup], and low values of $q$ are associated with brittle deformation, in which the stress concentration grows quickly with aggregate size. In contrast, cyanobacterial colonies have a soft plastic behavior that is less sensitive to large cracks, thus justifying the large value of $q$.
@@
We further tested the model with the aggregation experiments. A suspension of $C_{1}$ colonies with an initial distribution close to the single-cell size (Fig. 3A) was simulated. The model agrees well with the experimental results for the median colony size, accurately predicting a steady-state value when there is a balance between collisions and fragmentation (Fig. 3B). The best fit of the results provided a small value for the stickiness parameter ($\alpha_{1}=0.023$). The aggregation rate is linearly dependent on the stickiness and the total biovolume fraction (Eq. 2), thus an increase in these parameters ($\alpha_{1}$ and $\phi$) allows the suspension to reach the steady-state size faster. However, the steady-state size has low sensitivity to these parameters ($\alpha_{1}$ and $\phi$), as the fragmentation frequency has a higher-order dependence with colony size (Fig. 2F). The equal fragments hypothesis leads to a single-peaked size distribution, whose steady-state median increases with the total biovolume fraction. If an ideal erosion mechanism was assumed for the $C_{1}$ colonies, the peak would be fixed around the single-cell size (SI Appendix, Fig. S6). The experimental results for the $C_{1}$ colonies are in better agreement with the equal fragments hypothesis (Fig. 3C).
@@
Note that the ideal erosion and equal fragments mechanisms are idealized scenarios, and the real distribution is expected to be more dispersed. Furthermore, a fractal description of small colonies is limited, as a continuum function between biovolume and diameter of colonies may not hold. Moreover, the turbulent model used for fragmentation (Eq. 4) poses two limitations: (i) the theory was developed for homogeneous Gaussian turbulence [baebler2008modelling], which differs considerably from the boundary-layer turbulence generated in a cone-and-plate shear [grad2000spectral], and (ii) the colonies were larger than the Kolmogorov scale for the intense dissipation rate used in our fragmentation measurements ($\eta=20~{}\mu m$ for a dissipation rate of $\dot{\varepsilon}=5.8~{}{m^{/}s^{3}}$, SI Appendix), meaning that the colony size is comparable to the smallest eddies in the flow. Despite these limitations, the proposed two-category model provides a good approximation of the fragmentation process observed in our experiments. Simulations that take into account spatially inhomogeneous and non-Gaussian turbulence are demanding and lack an explicit analytic expression for $\kappa(l)$ that can be applied to a population model [babler2015numerical].
@@
Fig. 4: Kinetics of the fragmentation of colonies in field samples of Microcystis spp. at an intense dissipation rate ($\dot{\varepsilon}=5.8~{}{m^{/}s^{3}}$) and total biovolume fraction of $\phi=10^{-4}$. (A) Comparison of the fragmentation frequency as a function of colony size for the laboratory culture (Microcystis strain V163) and the field samples (Microcystis spp.). Error bars indicate the standard deviation. (B) Brightfield microscopy images of colonies in a Nigrosin-dyed medium (dark region) show evidence of a thick EPS layer (bright region) surrounding a field colony. (C) Initial size distribution of colonies in field samples as a function of the relative diameter. The size distribution had a bimodal shape, with small colonies (green) and large colonies (yellow). (D) The median diameter of the colonies in each subpopulation as a function of time. Bars indicate 25th and 75th percentiles. (E) After 1 hour of shear flow, the small colonies have aggregated slightly, while the large colonies keept their size distribution. (F) The fraction of small colonies remained nearly constant during the experiment.
Samples of Microcystis spp. were collected from Lake Gaasperplas - the Netherlands (Materials and Methods). A major difference in fragmentation behavior is observed between the laboratory cultures of Microcystis strain V163 and the field samples of Microcystis spp. While the laboratory cultures suffered intense fragmentation at a dissipation rate of $\dot{\varepsilon}=5.8$ ${m^{2}/s^{3}}$, the colonies collected from the field showed negligible fragmentation at this dissipation rate (Fig. 4A). The size distribution of field colonies also had a bimodal shape both initially and after 1 hour of flow (Figures 4C and E). The small colonies displayed an increase in size due to aggregation up to 2.5x times the initial median size (Fig. 4D), although the fraction of small colonies remained below $11\%$ even after a few hours of shear (Fig. 4F). The large colonies displayed only a small increase in size due to aggregation and quickly reached a steady-state size.
@@
The dimensionless ratio $Ag$ reveals the dominant mechanism for colony formation (cell division is dominant for $Ag\ll 1$, while aggregation is dominant for $Ag\gg 1$) and depends on the total biovolume fraction $\phi$ and dissipation rate $\dot{\varepsilon}$, as shown by the phase diagram in Figure 5. For a lake with a total biovolume fraction corresponding to an early bloom ($\phi=3\cdot 10^{-7}$, according to WHO alert level 1 [chorus2021toxic]), a low dissipation rate typical of wind-induced turbulence ($\dot{\varepsilon}\sim 10^{-7}~{}{m^{2}/s^{3}}$) [wuest2003small], a specific growth rate of $\mu\sim 0.008\,\text{h}^{-1}$ [huisman2004changes], a kinematic viscosity of $\nu=10^{-6}\text{m}^{2}\text{s}^{-1}$ and a stickiness of $\alpha\sim 0.02$, we estimate a ratio $Ag\sim 10^{-3}$. Under these conditions, the colony size increases due to cell division while the role of aggregation is negligible (region I in Fig. 5). In contrast, during a dense bloom event, the scum layer can reach a very high biovolume fraction of $\phi\sim 10^{-2}$ [wu2020recovery, wu2019effects], which leads to $Ag\sim 10^{2}$ (keeping all remaining parameters constant). Therefore, colony formation in dense scum layers is dominated by colony aggregation (region II in Fig. 5). The aggregation process can form larger colonies which will have an enhanced buoyancy, thus assisting in the persistence of the scum layer.
@@
Fragmentation limits the maximum colony size to a critical value $l^{*}$. Our experimental results (Fig. 5) indicate that aggregated colonies have a lower critical size than division-formed colonies for the same turbulence intensity. Conversely, a higher critical dissipation rate is needed for division-formed colonies ( $\dot{\varepsilon}_{div}$) than for aggregated colonies ($\dot{\varepsilon}_{agg}$) to achieve the same critical size $l^{*}=l_{m}$. This creates an interesting regime at high biovolume fractions and dissipation rates in the range $\dot{\varepsilon}_{agg}<\dot{\varepsilon}<\dot{\varepsilon}_{div}$, in which only division-formed colonies can survive the turbulent flow, even though aggregation is faster than division ($Ag\gg 1$) (region III in Fig. 5). Typical values of critical dissipation rate for our experiments are $\dot{\varepsilon}_{agg}=2.1\cdot 10^{-3}~{}{m^{2}/s^{3}}$ and $\dot{\varepsilon}_{div}=3.7~{}{m^{2}/s^{3}}$, assuming a maximum colony size of $l_{m}=50$. Under natural turbulent conditions (e.g. surface wind-mixing), the dissipation rate is normally below $\dot{\varepsilon}\sim 10^{-6}~{}{m^{/}s^{3}}$ [wuest2003small], which is insufficient to induce colony fragmentation. Artificial mixing systems can achieve levels of turbulence that are many orders of magnitude higher. In the core of bubble plumes, dissipation rates of $\dot{\varepsilon}\sim 10^{-4}-10^{-1}~{}{m^{2}/s^{3}}$ are reached [lai2019turbulent]. This value was sufficient to prevent aggregation of large colonies in our experiments but was not high enough to induce fragmentation of division-formed colonies. Moreover, colonies remain only a short time ($\sim 1$ min) in the bubble plume core, after which they are dispersed to weaker turbulent regions (dissipation rates of $\dot{\varepsilon}\sim 10^{-7}-10^{-4}~{}{m^{2}/s^{3}}$ just a few meters away from the bubble plume core). The intense hydrodynamic stress required to fragment division-formed colonies (region IV in Fig. 5) can only be achieved by mixers with fast-moving surfaces, such as the cone-and-plate setup used in our experiments ($\dot{\varepsilon}\sim 10^{-2}-10^{1}~{}{m^{2}/s^{3}}$) or propellers used in stirred tanks ($\dot{\varepsilon}\sim 10^{0}-10^{2}~{}{m^{2}/s^{3}}$) [laufhutte1987local, peter2006hydromechanical]. These mixing devices can provide intense dissipation rates in small fluid volumes, but their implementation across entire lakes would be excessively energy-consuming.
@@
Fig. 5: Phase diagram indicating the dominant colony formation mechanism as a function of the dissipation rate $\dot{\varepsilon}$ and the cyanobacterial abundance (expressed by the total biovolume fraction $\phi$). (I) Colonies grow only by cell division at low dissipation rates and total biovolume fractions. (II) As the biovolume fraction increases, aggregation enhances colony growth. (III) For moderate dissipation rates, aggregated colonies are fragmented, and only cell division can increase colony size. (IV) Fragmentation of colonies dominates at intense dissipation rates, irrespective of whether these colonies were formed by aggregation or cell division. Bars on the right side indicate typical values of dissipation rate. The horizontal arrowed line indicates the transition from the absence of a bloom to a dense scum formation under typical wind mixing, with the bullets indicating the total biovolume fraction for WHO alert level 1 [chorus2021toxic] (left) and for a typical scum layer [wu2020recovery] (right).
@@ Conclusions
We have demonstrated that flow-induced aggregated colonies of Microcystis form weaker structures than division-formed colonies, likely due to the weak EPS bonds between cells. Furthermore, we showed that colony aggregation is unlikely in lakes under early bloom conditions, as the collision frequency of colonies is too low. Our results, based on the mechanical properties of colonies, are corroborated by previous studies based on the sequencing of individual Microcystis colonies [perez2021single, smith2021individual]. These studies have identified that the genetic diversity within Microcystis colonies is much smaller than the diversity between Microcystis colonies, thus supporting the hypothesis that colony formation is dominated by cell division rather than by aggregation of colonies. Nonetheless, our results demonstrate that during scum formation in very dense blooms or phytoplankton patches, flow-induced aggregation likely plays a role and may quickly increase colony size. A summary of these outcomes is presented in the phase map illustrated in Figure 5.
From a broader perspective, our findings have relevance beyond predicting and mitigating cyanobacterial blooms. Since fluid flow affects the colonies and aggregates of many species [burd2009particle, drescher2013biofilm, rusconi2014bacterial, goldstein2015green, thomen2017bacterial, ros2024mixed], the experimental methods and theoretical framework presented here, along with the identified mechanisms, could serve as a foundation to better understand the relationship between mechanics and biological functions and traits of colonies under hydrodynamic stresses and their morphodynamics and lifestyle. Other examples include non-natural settings, such as algae or cyanobacteria cultured in bioreactors, where dense cell populations experience intense mixing [ducat2011engineering, parmar2011cyanobacteria, johnson2018photobioreactor]. The results presented here could serve as design guidelines for such applications.

### v2

@@ Introduction
To complement the mathematical model and to further investigate the mechanical origins of the resistance of division-formed colonies to intense shear, we performed rheology tests on the cyanobacterial colonies (Materials and Methods). Simple and oscillatory shear measurements revealed elasto-viscoplastic mechanics [balmforth2014yielding, bonn2017yield] of the concentrated aggregate with a yield stress of $\tau_{y}=4.3\pm 0.3\text{ Pa}$. The low volume fraction of cells in aggregate ($\sim 12$) indicates that the mechanical properties are mainly dictated by the EPS layer. Equating the average hydrodynamic shear stress to the yield stress in Eq. (11) and subsequently substituting the critical dissipation rate in Eq. (3) results in a critical colony size of $l^{*}_{2}=2.8$. The eroded fragments observed in the fragmentation experiments (Fig. 2E) are mostly smaller than this critical size. This suggests that large colonies will suffer erosion when the local hydrodynamic stress exceeds the yield stress of the EPS layer. Although the fracture behavior of colonies in suspension is expected to be different from a concentrated aggregate under shear rheology, the latter provides a reasonable and convenient estimate of the material properties of the colonies. Related to the critical size of the colony and their mechanical properties is the exponent $q_{2}$ (see equation 5). The best fit of the results in Fig. 2 provided $q_{2}=4.1$. This value lies in the upper limit of previous results for colloidal aggregates [flesch1999laminar, baebler2008modelling]. Note that for colloids, $q$ is related to the stress concentration around cracks in the fractal structure of the aggregate [zaccone2009breakup], and low values of $q$ are associated with brittle deformation, in which the stress concentration grows quickly with aggregate size. In contrast, cyanobacterial colonies have a soft plastic behavior that is less sensitive to large cracks, thus justifying the large value of $q$.
@@
We further tested the model with the aggregation experiments. A suspension of $C_{1}$ colonies with an initial distribution close to the single-cell size (Fig. 3A) was simulated with an equal fragments hypothesis. The model agrees well with the experimental results for the median colony size, accurately predicting a steady-state value when there is a balance between collisions and fragmentation (Fig. 3B). The best fit of the results provided a small value for the stickiness parameter ($\alpha_{1}=0.023$). The aggregation rate is linearly dependent on the stickiness and the total biovolume fraction (Eq. 2), thus an increase in these parameters ($\alpha_{1}$ and $\phi$) allows the suspension to reach the steady-state size faster. However, the steady-state size has low sensitivity to these parameters ($\alpha_{1}$ and $\phi$), as the fragmentation frequency has a higher-order dependence with colony size (Fig. 2F). The equal fragments hypothesis leads to a single-peaked size distribution, whose steady-state median increases with the total biovolume fraction. If an ideal erosion mechanism was assumed for the $C_{1}$ colonies, the peak would be fixed around the single-cell size (SI Appendix, Fig. S6). The experimental results for the $C_{1}$ colonies are in better agreement with the equal fragments hypothesis (Fig. 3C).
@@
Note that the ideal erosion and equal fragments mechanisms are idealized scenarios, and the real distribution is expected to be more dispersed. Furthermore, a fractal description of small colonies is limited, as a continuum function between biovolume and diameter of colonies may not hold. Moreover, the turbulence model used for fragmentation (Eq. 5) poses two limitations: (i) the theory was developed for homogeneous Gaussian turbulence [baebler2008modelling], which differs considerably from the boundary-layer turbulence generated in a cone-and-plate shear [grad2000spectral], and (ii) the colonies were larger than the Kolmogorov scale for the intense dissipation rate used in our fragmentation measurements ($\eta=20~{}\mu m$ for a dissipation rate of $\dot{\varepsilon}=5.8~{}{m^{2}/s^{3}}$, SI Appendix), meaning that the colony size is comparable to the smallest eddies in the flow. Despite these limitations, the proposed two-category model provides a good approximation of the fragmentation process observed in our experiments. Simulations that take into account spatially inhomogeneous and non-Gaussian turbulence are demanding and lack an explicit analytic expression for $\kappa(l)$ that can be applied to a population model [babler2015numerical].
@@
Fig. 4: Kinetics of the fragmentation of colonies in field samples of Microcystis spp. at an intense dissipation rate ($\dot{\varepsilon}=5.8~{}{m^{2}/s^{3}}$) and total biovolume fraction of $\phi=10^{-4}$. (A) Comparison of the fragmentation frequency as a function of colony size for the laboratory culture (Microcystis strain V163) and the field samples (Microcystis spp.). Error bars indicate the standard deviation. (B) Brightfield microscopy images of colonies in a Nigrosin-dyed medium (dark region) show evidence of a thick EPS layer (bright region) surrounding a field colony. (C) Initial size distribution of colonies in field samples as a function of the relative diameter. The size distribution had a bimodal shape, with small colonies (green) and large colonies (yellow). (D) The median diameter of the colonies in each subpopulation as a function of time. Bars indicate 25th and 75th percentiles. (E) After 1 hour of shear flow, the small colonies have aggregated slightly, while the large colonies keept their size distribution. (F) The biovolume fraction of small colonies remained nearly constant during the experiment.
Samples of Microcystis spp. were collected from Lake Gaasperplas - the Netherlands (Materials and Methods). A major difference in fragmentation behavior is observed between the laboratory cultures of Microcystis strain V163 and the field samples of Microcystis spp. While the laboratory cultures suffered intense fragmentation at a dissipation rate of $\dot{\varepsilon}=5.8$ ${m^{2}/s^{3}}$, the colonies collected from the field showed negligible fragmentation at this dissipation rate (Fig. 4A). The size distribution of field colonies also had a bimodal shape both initially and after 1 hour of flow (Figures 4C and E). The small colonies displayed an increase in size due to aggregation up to 2.5 times the initial median size (Fig. 4D), although the fraction of small colonies remained below $11\%$ even after a few hours of shear (Fig. 4F). The large colonies displayed only a small increase in size due to aggregation and quickly reached a steady-state size.
@@
The dimensionless ratio $Ag$ reveals the dominant mechanism for colony formation (cell division is dominant for $Ag\ll 1$, while aggregation is dominant for $Ag\gg 1$) and depends on the total biovolume fraction $\phi$ and dissipation rate $\dot{\varepsilon}$, as shown by the phase diagram in Figure 5. For a lake with a total biovolume fraction corresponding to an early bloom ($\phi=3\cdot 10^{-7}$, according to WHO alert level 1 [chorus2021toxic]), a low dissipation rate typical of wind-induced turbulence ($\dot{\varepsilon}\sim 10^{-7}~{}{m^{2}/s^{3}}$ at a depth of $\sim 1~{}\text{m}$) [wuest2003small], a specific growth rate of $\mu\sim 0.008\,\text{h}^{-1}$ [huisman2004changes], a kinematic viscosity of $\nu=10^{-6}\text{m}^{2}\text{s}^{-1}$ and a stickiness of $\alpha\sim 0.02$, we estimate a ratio $Ag\sim 10^{-3}$. Under these conditions, the colony size increases due to cell division while the role of aggregation is negligible (region I in Fig. 5). In contrast, during a dense bloom event, the scum layer can reach a very high biovolume fraction ($\phi\sim 10^{-2}$ [wu2020recovery, wu2019effects]) in the top layer (depth of $\sim 10^{-2}~{}\text{m}$), where concomitantly the dissipation rate is much higher ($\dot{\varepsilon}\sim 10^{-4}~{}{m^{2}/s^{3}}$ under weak wind turbulence [wang2016field]), which leads to $Ag\sim 10^{5}$ (keeping all remaining parameters constant). Although this regime (region II in Fig. 5) was not explored in our cone-and-plate experiments, recent mesocosm experiments found clear evidence for the aggregation of Microcystis colonies in surface scums, with rapid increase in colony size at rates exceeding cell division rates [wu2024dynamics]. Specifically, for a bulk biovolume fraction of $\phi\sim 10^{-4}$ and a dissipation rate of $\dot{\varepsilon}\sim 10^{-4}~{}{m^{2}/s^{3}}$ , that study measured a rate of increase in the average colony size of $\overline{l}^{-1}d\overline{l}/dt\sim 0.2~{}\text{h}^{-1}$, corresponding to a ratio of $Ag\sim 10^{1}$. These results lie well inside region II of Fig. 5. These mesocosm experiments confirm our model prediction that aggregation is the dominant mechanism for colony formation in region II.
@@
Fragmentation limits the maximum colony size to a critical value $l^{*}$. Our experimental results (Fig. 5) indicate that aggregated colonies have a lower critical size than division-formed colonies for the same turbulence intensity. Conversely, a higher critical dissipation rate is needed for division-formed colonies ( $\dot{\varepsilon}_{div}$) than for aggregated colonies ($\dot{\varepsilon}_{agg}$) to achieve the same critical size $l^{*}=l_{m}$. This creates an interesting regime at high biovolume fractions and dissipation rates in the range $\dot{\varepsilon}_{agg}<\dot{\varepsilon}<\dot{\varepsilon}_{div}$, in which only division-formed colonies can survive the turbulent flow, even though aggregation is faster than division ($Ag\gg 1$) (region III in Fig. 5). Typical values of critical dissipation rate for our experiments are $\dot{\varepsilon}_{agg}=2.1\cdot 10^{-3}~{}{m^{2}/s^{3}}$ and $\dot{\varepsilon}_{div}=3.7~{}{m^{2}/s^{3}}$, assuming a maximum colony size of $l_{m}=50$. Under natural turbulent conditions (e.g., surface wind-mixing), the dissipation rate ranges from $\dot{\varepsilon}\sim 10^{-8}~{}{m^{2}/s^{3}}$ [wuest2003small] to peaks of $\dot{\varepsilon}\sim 10^{-3}~{}{m^{2}/s^{3}}$ at strong winds (in the absence of wave breaking), which is still insufficient to induce colony fragmentation [wang2016field]. Artificial mixing systems can achieve levels of turbulence that are many orders of magnitude higher. In the core of bubble plumes, dissipation rates of $\dot{\varepsilon}\sim 10^{-4}-10^{-1}~{}{m^{2}/s^{3}}$ are reached [lai2019turbulent]. This value was sufficient to prevent aggregation of large colonies in our experiments but was not high enough to induce fragmentation of division-formed colonies. Moreover, colonies remain only a short time ($\sim 1$ min) in the bubble plume core, after which they are dispersed to weaker turbulent regions (dissipation rates of $\dot{\varepsilon}\sim 10^{-7}-10^{-4}~{}{m^{2}/s^{3}}$ just a few meters away from the bubble plume core). Nevertheless, bubble plumes can prevent the formation of a scum layer not only by fragmentation of newly formed weak aggregates, but also by providing sufficient turbulent mixing to disperse the buoyant colonies, thereby reducing the local biovolume fraction at the surface [visser2016artificial]. The intense hydrodynamic stress required to fragment division-formed colonies (region IV in Fig. 5) can only be achieved by mixers with fast-moving surfaces, such as the cone-and-plate setup used in our experiments ($\dot{\varepsilon}\sim 10^{-2}-10^{1}~{}{m^{2}/s^{3}}$) or propellers used in stirred tanks ($\dot{\varepsilon}\sim 10^{0}-10^{2}~{}{m^{2}/s^{3}}$) [laufhutte1987local, peter2006hydromechanical]. These mixing devices can provide intense dissipation rates in small fluid volumes, but their implementation across entire lakes would be excessively energy-consuming.
@@
Fig. 5: Phase diagram indicating the dominant colony formation mechanism as a function of the dissipation rate $\dot{\varepsilon}$ and the cyanobacterial abundance (expressed by the total biovolume fraction $\phi$). (I) Colonies grow only by cell division at low dissipation rates and total biovolume fractions. (II) As the biovolume fraction increases, aggregation enhances colony growth. (III) For moderate dissipation rates, aggregated colonies are fragmented, and only cell division can increase colony size. (IV) Fragmentation of colonies dominates at intense dissipation rates, irrespective of whether these colonies were formed by aggregation or cell division. Bars on the right side indicate typical values of dissipation rate observed for natural wind-mixing, bubble plumes in artificially mixed lakes and laboratory-scale setups such as cone-and-plate systems and stirred tanks. The dashed arrowed line indicates the transition from an early bloom (left bullet - WHO alert level 1 [chorus2021toxic]) to a dense surface scum (right bullet - typical scum biovolume fraction [wu2020recovery]) under typical wind mixing.
@@ Conclusions
We have demonstrated that flow-induced aggregated colonies of Microcystis form weaker structures than division-formed colonies, likely due to the weak EPS bonds between cells. Furthermore, we showed that colony aggregation is unlikely in lakes under early bloom conditions, as the collision frequency of colonies is too low. Our results, based on the mechanical properties of colonies, are corroborated by previous studies based on the sequencing of individual Microcystis colonies [perez2021single, smith2021individual]. These studies have identified that the genetic diversity within Microcystis colonies is much smaller than the diversity between Microcystis colonies, thus supporting the hypothesis that colony formation is dominated by cell division rather than by aggregation of colonies. Nonetheless, our model supports recent studies which demonstrated that, during scum formation in very dense blooms or phytoplankton patches, flow-induced aggregation likely plays a role and may quickly increase colony size [wu2019effects, wu2020recovery]. A summary of these outcomes is presented in the phase map illustrated in Figure 5.
From a broader perspective, the technological advance of combining microscopic imaging with rheometry has relevance beyond predicting and mitigating cyanobacterial blooms. Since fluid flow affects the colonies and aggregates of many species [burd2009particle, drescher2013biofilm, rusconi2014bacterial, goldstein2015green, thomen2017bacterial, ros2024mixed, song2022novel], the experimental methods and theoretical framework presented here, along with the identified mechanisms, could serve as a foundation to better understand the relationship between fluid mechanics and biological aggregates, while keeping a keen eye on the wide variation of biological functions and morphodynamics present in biological aggregates exposed to hydrodynamic stresses. An interesting example is a recent study on the fragmentation of marine snow consisting of diatoms and microplastics, which used similar methods as in our study [song2022novel, song2023deformation]. In contrast to the high mechanical resistance of Microcystis colonies, these marine snow aggregates displayed a low critical dissipation rate. Furthermore, their fragmentation was not mediated by an erosion mechanism, but by neck formation (i.e., deformation is higher in a small cross-sectional area in the middle of the aggregate). Other examples include non-natural settings, such as algae or cyanobacteria cultured in bioreactors, where dense cell aggregates experience intense mixing [ducat2011engineering, parmar2011cyanobacteria, johnson2018photobioreactor]. The methods and results presented here could serve as design guidelines for the study of these applications.